OpenStudy (anonymous):

3). To reduce the following rational expression to lowest terms, you will have to use factoring by grouping. Be sure to factor each numerator and denominator completely before dividing out any common factors. (Remember, factoring by grouping takes two steps.) (xy + 3 x + 4 y + 12)/((xy + 3 x + 9 y + 27))

OpenStudy (anonymous):

xy+3x+4y+12=x(y+3)+4(y+3)=(x+4)(y+3) xy+3x+9y+27=x(y+3)+9(y+3)=(x+9)(y+3) so we have (x+4)(y+3)/((x+9)(y+3)) and we will cancel y+3 \[(x+4)/(x+9)\]